# SHED: Shapley-Based Automated Dataset Refinement for Instruction Fine-Tuning

### v1

@@ Abstract
The pre-trained Large Language Models (LLMs) can be adapted for many downstream tasks and tailored to align with human preferences through fine-tuning. Recent studies have discovered that LLMs can achieve desirable performance with only a small amount of high-quality data, suggesting that a large amount of the data in these extensive datasets is redundant or even harmful. Identifying high-quality data from vast datasets to curate small yet effective datasets has emerged as a critical challenge. In this paper, we introduce SHED, an automated dataset refinement framework based on Shapley value for instruction fine-tuning. SHED eliminates the need for human intervention or the use of commercial LLMs. Moreover, the datasets curated through SHED exhibit transferability, indicating they can be reused across different LLMs with consistently high performance. We conduct extensive experiments to evaluate the datasets curated by SHED . The results demonstrate SHED ’s superiority over state-of-the-art methods across various tasks and LLMs; notably, datasets comprising only 10% of the original data selected by SHED achieve performance comparable to or surpassing that of the full datasets.
@@ 1 Introduction
The development of LLMs marks a major leap in machine learning, transforming how we approach natural language processing (NLP) and Artificial Intelligence (AI) research [1]. Models such as GPT-3 [2] and LLaMA [3] highlight the benefits of pre-training on large and diverse datasets, empowering these LLMs with a wealth of knowledge. Moreover, one of the pivotal strengths of LLMs lies in their adaptability to specific tasks through fine-tuning. Fine-tuning, a process that involves updating models on a task-specific dataset, enables the pre-trained model to acquire task-specific information. Furthermore, it facilitates the alignment of models to more accurately follow human instructions through fine-tuning on a dataset comprised of instructions paired with appropriate responses, which is known as instruction tuning.
@@
However, fine-tuning LLMs also raises challenges. A primary concern is that the noisy data or harmful instances in the fine-tuning dataset can significantly degrade the performance of pre-trained LLMs [4]. Many works have developed large and diverse datasets for fine-tuning purposes. While the datasets for fine-tuning have grown in size, recent research suggests that meticulously curated datasets of high quality, even if smaller in size, can be more effective in harnessing the full potential of LLMs [5, 6, 7]. Indiscriminately increasing the volume of data can lead to ineffective performance improvements and might even deteriorate model performance as a result of introducing noisy and harmful instances. In addition, for instruction tuning, the model has already learned the necessary knowledge in the pre-training stage. The dataset used in the fine-tuning stage merely aims to better align the models to follow human instructions, indicating this process does not necessitate extensive data [8]. Furthermore, fine-tuning LLMs on extensive datasets incurs significant computational costs. The necessity for considerable GPU resources presents a critical challenge [9]. Only those researchers and institutions equipped with sufficient computing resources are able to perform such tasks, limiting the broader applications and progress within the LLM community. Consequently, there is a pressing need to design a novel method for curating small and high-quality datasets that enable efficient fine-tuning.
@@
Previous efforts have employed various methods such as curation or generation through manual efforts or commercial LLMs [5, 10], identifying subsets from larger datasets via training dynamics or estimating marginal contributions [11, 12]. Most current methods for data selection neglect the potential influence that different combinations of samples can have on model performance. The Shapley value [13], introduced in cooperative game theory, provides a method for fairly evaluating the contribution of each participant by examining all possible combinations and their effects on the overall result. This principle has also been utilized in machine learning to assess the impact of individual data within a given dataset [14]. The Shapley value can serve as criteria to refine one or more large datasets to extract high-quality data points, enabling the curation of a smaller yet high-quality dataset. This method not only facilitates the selection of impactful data, but also considers the effectiveness of selected data combinations. Shapley value seems to be a promising toolkit for data selection. However, calculating the Shapley value for all the data samples in a dataset is computationally expensive, especially for large-scale fine-tuning datasets.
@@
Motivated by the aforementioned challenges, we present SHED , a Shapley-based automated dataset refinement framework for fine-tuning LLMs. The key intuition behind SHED is to perform Shapley value evaluations on a small portion of representative samples only, such that the computation complexity of the Shapley-based data refinement can be dramatically decreased. Specifically, as Figure 1 illustrates, SHED consists of three key components: model-agnostic clustering, proxy-based Shapley calculator, and optimization-aware sampling. Initially, the model-agnostic clustering groups embeddings of the original dataset, and then selects representative data samples as a proxy for each cluster based on the distance of embeddings to the cluster centroid. These proxy data instances are then evaluated by the proxy-based Shapley calculator, which employs an approximation method to efficiently calculate their Shapley values, focusing on task-specific objectives (e.g., accuracy and fairness). This method involves iteratively removing groups of instances from the proxy dataset and assessing the performance variation of the model to estimate the collective contribution of these instances, thereby streamlining the computation of Shapley values. The derived Shapley values of these proxy data instances are used as the quality score for their respective clusters. Finally, optimization-aware sampling selects data from clusters to compile a compact yet high-quality dataset, employing strategies that may favor clusters with higher-quality scores. SHED only computes Shapley values for the cluster representatives rather than each data, drastically boosting the efficiency of data refinement. Moreover, SHED offers a unified yet flexible framework, catering to various user needs by providing multiple options within each component. For example, the optimization objective for Shapley value measurement can be tailored to specific tasks (e.g., fairness). Our key contributions can be summarized as follows:
@@
  • We conducted extensive experiments on two benchmark datasets, i.e., MMLU and WizzardLLM, the results demonstrate that fine-tuning LLMs with small datasets curated by SHED yields performance comparable to, or even better than, using the original large datasets. Remarkably, we observed that the datasets curated by SHED exhibit transferability, achieving robust performance across various LLMs models even from different families. This amortizes the computational cost for data selection when applying the curated dataset to fine-tune various LLM models.
  • All the code associated with the collection of high-quality datasets curated by SHED can be found at SHED: Shapley-Based Automated Dataset Refinement.
@@ 2.1 Coreset Selection
Coreset selection plays a critical role in machine learning by targeting the selection of a representative subset from a larger dataset. Various coreset selection methods use unique criteria for choosing samples. Geometry-based approaches focus on the geometric properties of the data points, striving to retain geometrically significant samples that represent the overall data distribution [15, 16, 17, 18]. Uncertainty-based methods choose samples based on the uncertainty they present to the model, typically engaging samples that the model finds challenging to classify [19, 20, 21]. Decision-boundary-based methods select samples that are close to the decision boundary of the classifier, ensuring that the nuances of the classification boundary are well-represented in the selected subset [22, 23]. Gradient-matching approaches involve selecting a subset that yields similar gradient distributions as the entire dataset when used in training [24, 25]. Bilevel Optimization optimizes the coreset selection in a way that the selected subset maximizes certain performance metrics [26]. Submodularity-based approaches consider both diversity and information richness, striving for a balanced representation of the dataset [27].
@@ 2.2 Data Selection for Instruction Fine-tuning
Due to the superiority of instruction tuning in enhancing the performance of LLMs, many recent studies focus on selecting high-quality instruction tuning data. Based on methods, it can be divided into the following categories. Indicators-based methods define multiple metrics, such as instruction length and perplexity, to compute quality scores for each instruction instance [10, 28, 29, 30]. Training-based methods leverage the performance improvement through fine-tuning to score and select instruction data suited for fine-tuning [31, 32, 33, 34, 35, 36]. Some other methods employ commercial LLMs like ChatGPT to assess quality, complexity, and diversity of instructions for selection [7, 37, 38, 39, 40].
@@ 2.3 Limitations of Previous Work
Most existing methods for data selection overlook the impact of various data combinations on model performance. As Table 1 illustrates, datasets formed by combining high-quality data, which are merely based on the independent quality score of each individual data sample, do not necessarily enhance model performance effectively. The combination of different data can impact the final performance of fine-tuning. Furthermore, many works are task-specific, limiting their broader applicability.
@@ 3.1 Preliminary
Figure 2: Workflow of SHED: ① Clustering and determining proxy data; ② Calculating Shapley values as scores; and ③ Sampling based on scores.
@@ 4.1 Experimental Setup
Evaluation Settings. After obtaining the curated datasets using SHED and baseline methods, we fine-tune the pre-trained models using each curated subset, respectively. We apply the Low-Rank Adaptation (LoRA) for fine-tuning and set the default LoRA rank to 128. For all curated datasets, the instruction fine-tuning was conducted for 3 epochs. We evaluate the performance of fine-tuned models on MMLU and ARC-challenge tasks using the lm-evaluation-harness testing framework [47]. To better evaluate the human preferences of fine-tuned models, we adopt MT-Bench [48] in our experiments. All the experiments are conducted on two A100 GPUs.
@@ 4.2 Experiment Results
Evaluations on MT-Bench. We also use MT-Bench to evaluate the performance of datasets curated by SHED in terms of human preferences. Table 5 demonstrates that the dataset curated by SHED aligns well with human preferences, not only enhancing accuracy but also enabling the model to better understand and follow human instructions, generating answers that are more favorable to humans. The dataset constructed through the SHED-QWCS method, sampled from WizardLM, achieved a remarkable score of 5.24 on the MT-Bench.
Transferability Evaluation of Curated Datasets across Various Models. To evaluate the transferability of datasets curated by SHED , we first apply SHED to select data from the MMLU and WizardLM datasets based on LLaMA-7B. Then, we fine-tune LLaMA-13B, Vicuna-7B, and GPT-2 using the best-selected dataset curated by SHED and the baseline methods. As summarized in Table 6 and Table 7, datasets curated by SHED exhibit robust performance across various models, demonstrating their transferability and applicability across various tasks and even different model families. The strong transferability of the curated datasets indicates that SHED identifies generally high-quality data. In addition, the computational cost for data selection can be significantly amortized across various models.
@@
Impact of Number of Iterations on Proxy-based Shapley Calculator. The precision of Shapley value estimates increases with the number of iterations $k$, providing a more accurate measurement of each data sample’s contribution to the model performance. However, this increment also leads to a proportional rise in computational cost, leading to a contrasting relationship between computational efficiency and the accuracy of Shapley value estimations.
To seek the optimal number of iterations for Shapley value calculations, we analyzed the performance of datasets curated by SHED under varying iteration settings. The experiments are conducted with the LLaMA-7B model on the MMLU test set, following the experimental settings detailed in Section 4.1.
@@ 7 Limitations
This research, while presenting significant advancements, encounters certain limitations that merit attention for future work. Firstly, the proposed method’s reliance on clustering and proxy data may inadvertently diminish data diversity, potentially overlooking rare but critical samples. To address this, future research will investigate the adoption of novel clustering algorithms specifically designed to enhance the visibility and inclusion of rare samples. Secondly, the framework’s current objective focuses predominantly on model performance, which may inadvertently lead to model bias. This singular focus overlooks the equally important aspect of model fairness, which is crucial for ensuring that the models perform equitably across diverse groups. Recognizing this, our framework is designed to be extensible and objective-agnostic, laying the groundwork for incorporating additional criteria. In subsequent research, we plan to integrate considerations of model fairness alongside performance.
@@ 8 Ethics Statement
In this work, we present SHED, a generic data refinement framework for data refinement utilizing Shapley values, aimed at assembling a compact yet effective dataset to boost the efficiency of the fine-tuning process of LLMs. This study carefully avoids ethical issues beyond standard AI concerns, leveraging properly cited publicly available Internet text data. This approach ensures adherence to ethical data use standards, reflecting our commitment to responsible research practices in the AI field.
Appendix A Automatic Hyperparameter Setting
To enhance user convenience, SHED introduces an automatic mechanism for setting hyperparameters.
Based on Figure 3(c), the time per iteration when calculating Shapley values exhibits an approximate linear relationship with the number of clusters, and similarly, the total time for computing Shapley values closely aligns linearly with the number of iterations. Thus, the computation time $t$ for Shapley values can be modeled as $t=\theta kC$. SHED randomly samples 2000 instances from the dataset to calculate the Shapley value for one iteration, recording this to determine $\theta$. To optimize $k$ to be close to 10 and the number of clusters near $3\sqrt{\lvert D\rvert}$, an optimization problem is formulated in Eq. 3.
[…]

### v2

@@ Abstract
The pre-trained Large Language Models (LLMs) can be adapted for many downstream tasks and tailored to align with human preferences through fine-tuning. Recent studies have discovered that LLMs can achieve desirable performance with only a small amount of high-quality data, suggesting that a large portion of the data in these extensive datasets is redundant or even harmful. Identifying high-quality data from vast datasets to curate small yet effective datasets has emerged as a critical challenge. In this paper, we introduce SHED, an automated dataset refinement framework based on Shapley value for instruction fine-tuning. SHED eliminates the need for human intervention or the use of commercial LLMs. Moreover, the datasets curated through SHED exhibit transferability, indicating they can be reused across different LLMs with consistently high performance. We conduct extensive experiments to evaluate the datasets curated by SHED. The results demonstrate SHED’s superiority over state-of-the-art methods across various tasks and LLMs; notably, datasets comprising only 10% of the original data selected by SHED achieve performance comparable to or surpassing that of the full datasets.
@@ 1 Introduction
The development of LLMs marks a major leap in machine learning, transforming how we approach natural language processing (NLP) and artificial intelligence (AI) research [1, 2, 3, 4, 5]. LLMs such as GPT-3 [2], Mistral [6], and LLaMA/LLaMA2 [3, 4] highlight the benefits of pre-training on large and diverse mixtures of data corpora, empowering these LLMs with a wealth of knowledge[7, 8]. Moreover, one of the pivotal strengths of LLMs lies in their adaptability to specific tasks through fine-tuning. Fine-tuning, a process that involves adapting LLMs to one or multiple task-specific datasets, enables the pre-trained LLM to acquire task-specific information. Furthermore, it facilitates the alignment of LLMs to more accurately follow human instructions through fine-tuning on a dataset comprised of instructions paired with appropriate responses[9], which is known as instruction tuning.
@@
However, fine-tuning LLMs also raises challenges. A primary concern is that noisy data or harmful instances in the fine-tuning dataset can significantly degrade the performance of pre-trained LLMs [10]. While many works have developed large and diverse datasets for fine-tuning purposes, recent research suggests that meticulously curated datasets of high quality, even if smaller in size, can be more effective in harnessing the full potential of LLMs [11, 12, 13]. Indiscriminately increasing the volume of data can lead to ineffective performance improvements and might even deteriorate LLM performance due to the introduction of noisy and harmful instances. Additionally, for instruction tuning, the LLM has already learned the necessary knowledge in the pre-training stage. The dataset used in the fine-tuning stage merely aims to better align the LLM to follow human instructions, indicating that this process does not necessitate extensive data [14]. Furthermore, fine-tuning LLMs on extensive datasets incurs significant computational costs. The necessity for considerable GPU resources presents a critical challenge [15]. Only researchers and institutions equipped with sufficient computing resources can perform such tasks, limiting broader applications and progress within the LLM community. Consequently, there is a pressing need to design a novel method for curating small and high-quality datasets that enable efficient fine-tuning.
@@
Previous efforts have employed various methods such as curation or generation through manual efforts or commercial LLMs [11, 16], identifying subsets from larger datasets via training dynamics or estimating marginal contributions [17, 18]. Most current methods for data selection neglect the potential influence that different combinations of samples can have on model performance. The Shapley value [19], introduced in cooperative game theory, provides a method for fairly evaluating the contribution of each participant by examining all possible combinations and their effects on the overall result. This principle has also been utilized in machine learning to assess the impact of individual data points within a given dataset [20]. The Shapley value can serve as a criterion to refine one or more large datasets to extract high-quality data points, enabling the curation of a smaller yet high-quality dataset. This method not only facilitates the selection of impactful data but also considers the effectiveness of selected data combinations. The Shapley value seems to be a promising tool for data selection. However, calculating the Shapley value for all the data samples in a dataset is computationally expensive, especially for large-scale fine-tuning datasets.
Motivated by the aforementioned challenges, we present SHED, a Shapley-based automated dataset refinement framework for fine-tuning LLMs. The key intuition behind SHED is to perform Shapley value evaluations on a small portion of representative samples only, thereby dramatically decreasing the computational complexity of Shapley-based data refinement.
@@
Specifically, as Figure 1 illustrates, SHED consists of three key components: (1) model-agnostic clustering, (2) proxy-based Shapley calculator, and (3) optimization-aware sampling. Initially, the model-agnostic clustering groups embeddings of the original dataset and then selects representative data samples as a proxy for each cluster based on the distance of embeddings to the cluster centroid. These proxy data instances are then evaluated by the proxy-based Shapley calculator, which employs an approximation method to efficiently calculate their Shapley values, focusing on task-specific objectives (e.g., accuracy and fairness). This method involves iteratively removing groups of instances from the proxy dataset and assessing the performance variation of the model to estimate the collective contribution of these instances, thereby streamlining the computation of Shapley values. The derived Shapley values of these proxy data instances are used as the quality score for their respective clusters. Finally, optimization-aware sampling selects data from clusters to compile a compact yet high-quality dataset, employing strategies that may favor clusters with higher-quality scores.
SHED only computes Shapley values for the cluster representatives rather than each data point, drastically boosting the efficiency of data refinement. Furthermore, Yang et al. (2022) observed that hyperparameters tuned on smaller models can be effectively transferred to larger models, significantly reducing tuning costs while maintaining performance [21]. We observed a similar phenomenon: datasets curated by SHED exhibit strong transferability, performing robustly across LLMs of various sizes and families. This suggests that smaller LLMs can be used for data selection, reducing computational costs. The selected datasets can be used to fine-tune larger LLMs and reused in multiple tasks to further amortize costs. Moreover, SHED offers a unified yet flexible framework, catering to various user needs by providing multiple options within each component. For example, the optimization objective for Shapley value measurement can be tailored to specific tasks (e.g., fairness).
Our key contributions can be summarized as follows:
@@
  • We conducted extensive experiments on two benchmark datasets, i.e., MMLU and WizzardLM, the results demonstrate that fine-tuning LLMs with small datasets curated by SHED yields performance comparable to, or even better than, using the original large datasets. Notably, datasets curated by SHED exhibit strong transferability, achieving robust performance across various LLMs of different sizes and families. This indicates that smaller models can employed to greatly lower computational expenses for data selection, and the selected dataset can be used to fine-tune larger models and reused across multiple tasks to further distribute the costs.
  • Code associated with the collection of high-quality datasets curated by SHED can be found at SHED: Shapley-Based Automated Dataset Refinement.
@@ 2.1 Coreset Selection
Coreset selection plays a critical role in machine learning by targeting the selection of a representative subset from a larger dataset. Various coreset selection methods use unique criteria for choosing samples. Geometry-based approaches focus on the geometric properties of the data points, striving to retain geometrically significant samples that represent the overall data distribution [22, 23, 24, 25]. Uncertainty-based methods choose samples based on the uncertainty they present to the model, typically engaging samples that the model finds challenging to classify [26, 27, 28]. Decision-boundary-based methods select samples that are close to the decision boundary of the classifier, ensuring that the nuances of the classification boundary are well-represented in the selected subset [29, 30]. Gradient-matching approaches involve selecting a subset that yields similar gradient distributions as the entire dataset when used in training [31, 32]. Bilevel Optimization optimizes the coreset selection in a way that the selected subset maximizes certain performance metrics [33]. Dataset Selection with Datamodels using datamodels to approximate how the learning algorithm utilizes different subsets of training data to minimize target task loss.[34, 35] Submodularity-based approaches consider both diversity and information richness, striving for a balanced representation of the dataset [36].
@@ 2.2 Data Selection for Instruction Fine-tuning
Due to the superiority of instruction fine-tuning in enhancing the performance of LLMs, many recent studies focus on selecting high-quality instruction fine-tuning data. Based on methods, it can be divided into the following categories. Indicators-based methods define multiple metrics, such as instruction length and perplexity, to compute quality scores for each instruction instance [16, 37, 38, 39]. Training-based methods leverage the performance improvement through fine-tuning to score and select instruction data suited for fine-tuning [18, 40, 41, 42, 43, 44, 45]. Some other methods employ commercial LLMs like ChatGPT to assess quality, complexity, and diversity of instructions for selection [13, 46, 47, 48, 49].
@@ 2.3 Limitations of Previous Work
Most existing methods for data selection overlook the impact of various data combinations on model performance. As Table 1 illustrates, datasets formed by combining high-quality data, which are merely based on the independent quality score of each individual data sample, do not necessarily enhance model performance effectively. The combination of different data can impact the final performance of fine-tuning.
Although TS-DSHAPLEY [18] also utilized Shapley value for data selection, SHED offers several distinct advantages. SHED computes Shapley values only for proxy data of clusters rather than each individual data point, dramatically reducing computational overhead compared to TS-DSHAPLEY. SHED employs model-agnostic clustering, enhancing the transferability of curated datasets across different language models and model families. Moreover, SHED considers data diversity and can be customized for various optimization objectives, while TS-DSHAPLEY primarily focuses on predictive accuracy.
Many other existing works are also task-specific, limiting their applicability. In contrast, SHED offers a unified and flexible framework, adaptable to various instructional tuning tasks, making it more widely applicable.
@@ 3.2 Design of SHED
Figure 2: Workflow of SHED: ① Clustering and determining proxy data; ② Calculating Shapley values as scores; ③ Sampling based on scores; and ④ Forming the selected dataset.
@@ 4.1 Experimental Setup
Evaluation Settings. After obtaining the curated datasets using SHED and baseline methods, we fine-tune the pre-trained models using each curated subset, respectively. We apply the Low-Rank Adaptation (LoRA), which is a flexible and efficient tool, for fine-tuning and set the default LoRA rank to 128 [56, 57]. For all curated datasets, the instruction fine-tuning was conducted for 3 epochs. Notably, we use the same hyperparameters in fine-tuning across all methods to ensure a fair comparison, aiming to isolate the impact of data selection on model performance. We evaluate the performance of fine-tuned models on MMLU and ARC-challenge tasks using the lm-evaluation-harness testing framework [58]. To better evaluate the human preferences of fine-tuned models, we adopt MT-Bench [59] in our experiments. All the experiments are conducted on two A100 GPUs, each with 80GB of memory.
@@ 4.2 Experiment Results
Evaluations on MT-Bench. We use MT-Bench to evaluate the performance of datasets curated by SHED in terms of human preferences. Table 5 demonstrates that the datasets curated by SHED align well with human preferences, not only enhancing accuracy but also enabling the model to better understand and follow human instructions, generating answers that are more favorable to humans. The dataset constructed through the SHED-QWCS method, sampled from WizardLM, achieved a remarkable score of 5.24 on the MT-Bench. The results presented in Table 5 represent the average of five independent runs.
Transferability Evaluation of Curated Datasets across Various Models. To evaluate the transferability of datasets curated by SHED, we first apply SHED to select data from the MMLU and WizardLM datasets based on LLaMA-7B. Then, we fine-tune LLaMA-13B, Vicuna-7B, and GPT-2 using the best-selected dataset curated by SHED and the baseline methods. As summarized in Table 6 and Table 7, datasets curated by SHED exhibit robust performance across various models, demonstrating their transferability and applicability across various tasks and even different model families. The strong transferability of the curated datasets indicates that SHED identifies generally high-quality data. The computational cost for data selection can be significantly amortized across various models. In addition, the datasets selected by LLaMA-7B also achieve good performance when fine-tuning the larger model LLaMA-13B. This indicates that we can utilize smaller models to select data, thereby significantly reducing the computational cost of data selection.
@@
Impact of Number of Iterations on Proxy-based Shapley Calculator. The precision of Shapley value estimates increases with the number of iterations $k$, providing a more accurate measurement of each data sample’s contribution to the model performance. However, this increment also leads to a proportional rise in computational cost, leading to a contrasting relationship between computational efficiency and the accuracy of Shapley value estimations. To seek the optimal number of iterations for Shapley value calculations, we analyzed the performance of datasets curated by SHED under varying iteration settings. The experiments are conducted with the LLaMA-7B model on the MMLU test set, following the experimental settings detailed in Section 4.1.
@@ 7 Limitations
This research, while presenting significant advancements, encounters certain limitations that merit attention for future work. Firstly, the method’s reliance on sufficiently representative embeddings may limit its applicability in real-world scenarios where such embeddings are unavailable or inadequate. Future work will explore ways to reduce this dependency for broader applicability. Secondly, the use of clustering and proxy data may overlook rare but important samples. Future research will focus on improving clustering methods to better capture these samples. Additionally, the framework’s current objective focuses predominantly on model performance, which may inadvertently lead to model bias. This singular focus overlooks the equally important aspect of model fairness, crucial for ensuring that models perform equitably across diverse groups. Recognizing this, our framework is designed to be extensible and objective-agnostic, laying the groundwork for incorporating additional criteria. In subsequent research, we plan to integrate considerations of model fairness alongside performance.
@@ 8 Ethics Statement
In this work, we present SHED, a generic data refinement framework utilizing Shapley values, aimed at assembling a compact yet effective dataset to boost the efficiency of the fine-tuning process of LLMs. This study carefully avoids ethical issues beyond standard AI concerns, leveraging properly cited publicly available Internet text data. This approach ensures adherence to ethical data use standards, reflecting our commitment to responsible research practices in the AI field.
Acknowledgements
We thank the anonymous reviewers for their valuable insights and recommendations, which have greatly improved our work. This research has been graciously funded by NSF 2431611.
Appendix
Appendix A Hyperparameter Settings and Experimental Configuration
In our experiments, we employed the following hyperparameters: the number of training epochs was set to 3, the batch size was 128, the LoRA rank (lora_r) was 128, and the LoRA alpha (lora_alpha) was 256. For clustering, when the number of clusters (C) was 3000, the number of samples removed per group (n) was 60; when testing the impact of different C values on performance, $n=\frac{C}{50}$. The number of iterations for Shapley value calculation (k) was 10, and the learning rate was $3\times 10^{-4}$. Data preprocessing involved using the sentence-transformers/all-MiniLM-L6-v2 model to generate semantically meaningful embeddings for each sample in the original dataset, followed by applying the k-means algorithm to cluster these embeddings.